i need your help in this
a professor of statistics has two boxes of computer disks box c1 contains seven vebratim disks and 3 control data disks and box c2 contains 2 verbatim disks and 8 control data disks, she selects a box at random with probabilities p(c1)2/3 , and p(c2)=1/3 because their respective location, A disk is then selected at random and the event c occurs if it is from control data , using an equaly likely assumption for each disk in the selected box , compute p(c1|cc) abd p(c2|c)

Answer by stanbon(48568)

a professor of statistics has two boxes of computer disks box c1 contains seven vebratim disks and 3 control data disks and box c2 contains 2 verbatim disks and 8 control data disks, she selects a box at random with probabilities p(c1)2/3 , and p(c2)=1/3 because their respective location, A disk is then selected at random and the event c occurs if it is from control data , using an equaly likely assumption for each disk in the selected box , compute p(c1|cc) abd p(c2|c)
------------------------
Draw a contingency table with rows for c1 and c2 and columns for v's and c's
Calculate the row and the column totoals.
--------------------------
P(c1|cc) Did you mean P(c1|c)?
If so P(c1|c)=[P(c1 and c)]/[P(c)}= [(3/20)]/[(11/20)]=3/11
----------
P(c2|c)=[P(c2 and c)]/P(c) = [(8/20)]/[(11/20)]=8/11
------------
Cheers,
Stan H.